OpenStudy (anonymous):

For the series n=1 to infinity, 8/n^3, use the inequalities shown with n=10. Then, use the midpoint of the interval found to approximate the sum of the series.

OpenStudy (anonymous):

\[s_{n}+\int\limits_{n+1}^{\infty} f(x)dx \le S \le s_{n}+ \int\limits_{n}^{\infty} f(x)dx\]

OpenStudy (anonymous):

a)\[? \le \sum_{n=1}^{\infty} \frac{ 8 }{ n^{3} }\le?\]

OpenStudy (anonymous):

@rational

OpenStudy (rational):

Yeah start by finding \(s_n\)

OpenStudy (anonymous):

It isn't a geometric or telescopic series right?

OpenStudy (rational):

use wolfram to find \(s_{10}\)

OpenStudy (rational):

\[s_{10} = \sum\limits_{n=1}^{10} \dfrac{8}{n^3} = ?\]

OpenStudy (anonymous):

9.580

OpenStudy (rational):

Yes so we have \[\color{red}{9.580}+\int\limits_{10+1}^{\infty} f(x)dx \le S \le \color{red}{9.580}+ \int\limits_{10}^{\infty} f(x)dx\]

OpenStudy (rational):

plugin f(x) = 8/x^3 and evaluate the definite integrals above

OpenStudy (anonymous):

Alright, thanks!

OpenStudy (rational):

yw